OpenStudy (tayfishb):

find all critical points of the fxn Fx=x^3-9/2x^2-54x+2

myininaya (myininaya):

did you mean f(x)=(x^3-9)/(2x^2-54x+2) or is it f(x)=x^3-9/2x^2-54x+2?

myininaya (myininaya):

Whatever the function, the first step is to find the derivative

OpenStudy (tayfishb):

\[f(x)=x ^{3}-\frac{ 9 }{ 2 }x ^{2}-54x\]

myininaya (myininaya):

oh that is much easr that what I have imagined

myininaya (myininaya):

do you know power rule?

myininaya (myininaya):

\[(x^n)' = n \cdot x^{n-1}\]

OpenStudy (tayfishb):

x^n=nx^(n-1)

myininaya (myininaya):

yep apply that to each term

myininaya (myininaya):

\[(x^3)'=?\]

OpenStudy (tayfishb):

3x^2

myininaya (myininaya):

\[(-\frac{9}{2}x^{2})' =\frac{-9}{2}(x^2)'=?\]

OpenStudy (tayfishb):

-9x???

myininaya (myininaya):

yep

myininaya (myininaya):

and finally \[(-54x)'=?\]

OpenStudy (tayfishb):

-54

myininaya (myininaya):

right

myininaya (myininaya):

so put the three terms together and you are done

myininaya (myininaya):

\[f(x)=x^3-\frac{9}{2}x^2-54x \\ f'(x)=(x^3)'-\frac{9}{2}(x^2)'-54(x)' \\ f'(x)=3x^2-\frac{9}{2}(2x)-54(1) \\ f'(x)=3x^2-9x-54\]

myininaya (myininaya):

well done with the derivative

myininaya (myininaya):

you still need to find the critial numbers :p

myininaya (myininaya):

you can do that by replacing f'(x) with 0

OpenStudy (tayfishb):

ok so whats the next move?

myininaya (myininaya):

\[0=3x^2-9x-54\] and solve

OpenStudy (tayfishb):

solve how exactly?

myininaya (myininaya):

it is a quadratic equation... you can try factoring or using the quadratic formula...

OpenStudy (tayfishb):

\[3(x ^{2}-3x-18)\]

myininaya (myininaya):

ok do you know how to factor x^2-3x-18?

OpenStudy (tayfishb):

not exactly. there arent any common terms

myininaya (myininaya):

do you know how to find two numbers that when you multiply them you get -18 but add them and get -3?

OpenStudy (tayfishb):

-6 and 3

myininaya (myininaya):

so the factored form of x^2-3x-18 is (x-6)(x+3)

OpenStudy (tayfishb):

oh my goodness sorry i should have realized that my brain is so fried so it would be 6 and 3?

myininaya (myininaya):

6 and -3

OpenStudy (tayfishb):

-3

OpenStudy (tayfishb):

thank you so much

myininaya (myininaya):

np